Add
(6x^2 - xy + y^2) + (-x^2 - 5xy + 7y^2)

Answer by joecbaseball(37)

The best way to do this is to line up your equations with “like” variable terms grouped together over each other. In this problem it is easy, since each of the three terms in the first equation is the same as the three terms in the second equation:
You can only add “like” terms. For instance, x terms can only be added to x terms. y^2 terms can only be added to y^2 terms, etc… Just like dogs cannot be added to cats if you’re trying to figure out how many dogs you have.
So….
6x^2 - xy + y^2
+
-x^2 - 5xy + 7y^2
Add the terms going down:
For the x^2 terms, you have 6 + (-1) = 5
For the xy terms, you have (-1) + (-5) = -6
For the y^2 terms, you have 1 + 7 = 8
Put them all together, and you have for an answer 5x^2 – 6xy + 8y^2.
Good luck !